Please help me with the problem below:
Give three examples of the equations of parabolas that have (0,5) as the y-axis intercept

Answer by Alan3354(69443)

Give three examples of the equations of parabolas that have (0,5) as the y-axis intercept
----------
y = x^2 + 5
y = x^2 + x + 5
y = x^2 - x + 5
-----------
y = ax^2 + bx + 5, for any values of a & b